OpenStudy (anonymous):

Physics Help!!! Your car tire has a diameter of 3m and is rotating at 4 rev/s, when suddenly Bob slams his foot down on the accelerator. After traveling 400 m, the tire's rotation has increased to 10 rev/s. What was the tire's angular acceleration?

OpenStudy (anonymous):

I know the formula for angular acceleration is this: \[\alpha=\frac{ d \omega }{ dt }\]

OpenStudy (anonymous):

Well, try this formula: w^2=w_0^2+2*al*th you can find th first by finding how many revolution the tire made in 400 m (if it's diameter is 3 m). Then you can substitute this to the first formula for al. Hope this help :D read: w=\omega; al=\alpha; th=\theta; sorry for this :)

OpenStudy (anonymous):

So the formula is: \[w^2=w_0^2+2*al*th\] What's the theta? I mean, what would I plug in for the theta. So, the alpha would be equal to 40... Is that right?

OpenStudy (anonymous):

According to the unit used in the problem, you can substitute theta with number of revolution (I stated this before, right?).